OpenStudy (dls):

Calculate the wave length of a wave if its frequency is 15 mega cycles per second and its velocity is 3*10^8m/s

OpenStudy (dls):

\[\LARGE \lambda=\frac{3 \times 10^{8}}{15 \times 10^6}\] this?

OpenStudy (dls):

@yrelhan4 @Mashy

OpenStudy (dls):

15 mega cycles means 15 MHz?

OpenStudy (dls):

@experimentX

OpenStudy (yrelhan4):

megacycles/sec matlab Mhz.

OpenStudy (experimentx):

looks okay ...

OpenStudy (dls):

but I multiplied by 10^3 in exam :/

OpenStudy (yrelhan4):

ohho. tune Mhz hi likha tha. mere laga Hz likha hai. :P koi na. pta hai. ye imp hai.

OpenStudy (experimentx):

Kilo herz = 3 mega = 6 ... you should know these basic stuff ... from memory unit in computers.

OpenStudy (dls):

so megacycles=megahertz=10^hertz=10^6 cycles?

OpenStudy (experimentx):

*hertz

OpenStudy (experimentx):

yes yes ..

OpenStudy (dls):

dam..okay..